OpenStudy (anonymous):

Convert each rate using dimensional analysis. Round to the nearest tenth, if necessary. 45 mi/h = ___ ft/s

OpenStudy (imstuck):

What is dimensional analysis? I can do this, but I'm not sure dimensional analysis is!

OpenStudy (anonymous):

1 mi = 1760 yards =1760 * 3 ft 1hr = 3600 sec

OpenStudy (anonymous):

Thank you, @matricked And yeah, thats the part where I got stuck on too @IMStuck

OpenStudy (anonymous):

hop eu can do further...

OpenStudy (anonymous):

45 mi/h= 45mi/(1hr) = 45 *1760*3 ft/(3600 s)

OpenStudy (anonymous):

hhmm...I must be going wrong somewhere as the answer I got wasn't on the multiple choice

OpenStudy (imstuck):

I do it like this:

OpenStudy (imstuck):